# Through the Looking Glass into the Dark Dimension: Searching for Bulk Black Hole Dark Matter with Microlensing of $\bm{X}$-ray Pulsars

### v1

@@ I Introduction
Primordial black holes (PBHs), presumably formed through the collapse of sufficiently sizable overdensities originated by an enhancement in the comoving curvature perturbation power spectrum at small scales during inflation Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx , may be one of the most interesting candidates of dark matter Carr:2020xqk ; Green:2020jor ; Villanueva-Domingo:2021spv . These intriguing dark matter candidates are particular important in the dark dimension scenario, because 5D PBH are bigger, colder, and longer-lived than usual 4D PBHs of the same mass $M$ Anchordoqui:2022txe ; Anchordoqui:2022tgp ; Anchordoqui:2024dxu . [^1] The dark dimension scenario provides a specific realization of the dynamical dark matter framework Dienes:2011ja with other interesting dark matter candidates which have been discussed in Gonzalo:2022jac ; Anchordoqui:2023tln ; Law-Smith:2023czn ; Obied:2023clp . Furthermore, we have recently shown that a back reaction of Hawking evaporation process is to kick 5D Schwarzschild PBHs out of the brane Anchordoqui:2024jkn . As a consequence, 5D PBHs could make all cosmological dark matter if $10^{11}\lesssim M/{\rm g}\lesssim 10^{21}$, i.e., well down into the asteroid-mass window. In this paper we investigate a potential signal that may allow us to unmask these as yet evasive 5D PBHs.
@@ III.1 Microlensing of Bulk Black Holes
and where $\Gamma(x)$ is the Gamma function Tangherlini:1963bw ; Myers:1986un ; Argyres:1998qn . The $d$-dimensional black hole behaves like a thermodynamic system Hawking:1976de , with temperature $T_{H}\sim(d-3)/(4\pi r_{s})$ and entropy $S=(4\pi Mr_{s})/(d-2)$ Anchordoqui:2001cg . If the black hole is localized on the brane, Hawking evaporation Hawking:1974rv ; Hawking:1975vcx proceeds dominantly through emission of Standard Model fields Emparan:2000rs . However, we have recently shown that the recoil effect due to graviton emission imparts the black hole a relative kick velocity with respect to the brane, allowing Schwarzschild black holes to escape into the bulk Anchordoqui:2024jkn . Because the escape from the brane is almost instantaneous 5D Schwarzschild black holes evaporate almost entirely into gravitons in the bulk, and so can evade constraints from the non-observation of Hawking radiation in: (i) the extragalactic $\gamma$-ray background Carr:2009jm , (ii) the cosmic microwave background Clark:2016nst , (iii) the 511 keV $\gamma$-ray line DeRocco:2019fjq ; Laha:2019ssq ; Dasgupta:2019cae ; Keith:2021guq , (iv) the EDGES 21-cm signal Mittal:2021egv , and (v) the MeV Galactic diffuse emission Laha:2020ivk ; Berteaud:2022tws ; Korwar:2023kpy .
@@
Figure 2: Compilation of projected sensitivities on the PBH dark matter fraction $f_{\rm PBH}$ as a function of $M$ for future $X$-ray telescopes, assuming a monochromatic mass function and including wave optics effects. The sensitivity curves have been taken from Bai:2018bej ; Tamta:2024pow . For comparison, the shaded region shows contraints on the 4D $f_{\rm PBH}$ from the non-observation of Hawking radiation in: (i) the extragalactic $\gamma$-ray background Carr:2009jm , (ii) the cosmic microwave background Clark:2016nst , (iii) the 511 keV $\gamma$-ray line DeRocco:2019fjq ; Laha:2019ssq ; Dasgupta:2019cae ; Keith:2021guq , (iv) the EDGES 21-cm signal Mittal:2021egv , and (v) the MeV Galactic diffuse emission Laha:2020ivk ; Berteaud:2022tws ; Korwar:2023kpy . .
@@ III.2 Microlensing of Near-Extremal Black Holes
We can see that the effect of $Q$ is to slightly reduce the second order correction to the deflection angle. Therefore, we can conclude that the sensitivity of future $X$-ray microlensing experiments to near-extremal black holes is actually shown in Fig. 2.

### v2

@@ I Introduction
Primordial black holes (PBHs), presumably formed through the collapse of sufficiently sizable overdensities originated by an enhancement in the comoving curvature perturbation power spectrum at small scales during inflation Zeldovich:1967lct ; Hawking:1971ei ; Carr:1974nx , may be one of the most interesting candidates of dark matter Carr:2020xqk ; Green:2020jor ; Villanueva-Domingo:2021spv . These intriguing dark matter candidates are particular important in the dark dimension scenario, because 5D PBH are bigger, colder, and longer-lived than usual 4D PBHs of the same mass $M$ Anchordoqui:2022txe ; Anchordoqui:2022tgp ; Anchordoqui:2024dxu . [^1] The dark dimension scenario provides a specific realization of the dynamical dark matter framework Dienes:2011ja with other interesting dark matter candidates which have been discussed in Gonzalo:2022jac ; Anchordoqui:2023tln ; Law-Smith:2023czn ; Obied:2023clp ; Gendler:2024gdo . Furthermore, we have recently shown that a back reaction of Hawking evaporation process is to kick 5D Schwarzschild PBHs out of the brane Anchordoqui:2024jkn . As a consequence, 5D PBHs could make all cosmological dark matter if $10^{11}\lesssim M/{\rm g}\lesssim 10^{21}$, i.e., well down into the asteroid-mass window. In this paper we investigate a potential signal that may allow us to unmask these as yet evasive 5D PBHs.
@@ III.1 Microlensing of Bulk Black Holes
and where $\Gamma(x)$ is the Gamma function Tangherlini:1963bw ; Myers:1986un ; Argyres:1998qn . The $d$-dimensional black hole behaves like a thermodynamic system Hawking:1976de , with temperature $T_{H}\sim(d-3)/(4\pi r_{s})$ and entropy $S=(4\pi Mr_{s})/(d-2)$ Anchordoqui:2001cg . If the black hole is localized on the brane, Hawking evaporation Hawking:1974rv ; Hawking:1975vcx proceeds dominantly through emission of Standard Model fields Emparan:2000rs . However, we have recently shown that the recoil effect due to graviton emission imparts the black hole a relative kick velocity with respect to the brane, allowing Schwarzschild black holes to escape into the bulk Anchordoqui:2024jkn . Because the escape from the brane is almost instantaneous 5D Schwarzschild black holes evaporate almost entirely into gravitons in the bulk, and so can evade constraints from the non-observation of Hawking radiation in: (i) the extragalactic $\gamma$-ray background Carr:2009jm , (ii) the cosmic microwave background Clark:2016nst , (iii) the 511 keV $\gamma$-ray line DeRocco:2019fjq ; Laha:2019ssq ; Dasgupta:2019cae ; Keith:2021guq , (iv) the EDGES 21-cm signal Mittal:2021egv ; Saha:2021pqf , (v) Lyman-$\alpha$ forest Saha:2024ies , and (vi) the MeV Galactic diffuse emission Laha:2020ivk ; Berteaud:2022tws ; Korwar:2023kpy .
@@
Figure 2: Compilation of projected sensitivities on the PBH dark matter fraction $f_{\rm PBH}$ as a function of $M$ for future $X$-ray telescopes, assuming a monochromatic mass function and including wave optics effects. The sensitivity curves have been taken from Bai:2018bej ; Tamta:2024pow . For comparison, the shaded region shows contraints on the 4D $f_{\rm PBH}$ from the non-observation of Hawking radiation in: (i) the extragalactic $\gamma$-ray background Carr:2009jm , (ii) the cosmic microwave background Clark:2016nst , (iii) the 511 keV $\gamma$-ray line DeRocco:2019fjq ; Laha:2019ssq ; Dasgupta:2019cae ; Keith:2021guq , (iv) the EDGES 21-cm signal Mittal:2021egv ; Saha:2021pqf , (v) Lyman-$\alpha$ forest Saha:2024ies , and (vi) the MeV Galactic diffuse emission Laha:2020ivk ; Berteaud:2022tws ; Korwar:2023kpy . .
@@ III.2 Microlensing of Near-Extremal Black Holes
We can see that the effect of $Q$ is to slightly reduce the second order correction to the deflection angle. Therefore, we can conclude that the sensitivity of future $X$-ray microforlensing experiments to near-extremal black holes is actually shown in Fig. 2.
In summary, evaporation constraints on $f_{\rm PBH}$ can be substantially altered when moving away from the Schwarzschild picture. (Other PBH models where Hawking radiation can be slowed down have been recently discussed in Calza:2024fzo ; Calza:2024xdh .) Microlensing experiments of $X$-ray pulsars are well positioned to uncover these models.